Find three consecutive integers such that four times the sum of all three is 2 times the product of the larger two.

Answer by ewatrrr(24785)


Hi,
Let x, (x+1) and (x+2) represent the three consecutive integers
Write as we Read***
4[x + (x+1) + (x+2)]=2(x+1)(x+2)
solving for s
4x + 4x + 4 + 4x + 8 = 2[x^2 + 3x + 2]
12x + 12 = 2x^2 + 6x + 4
2x^2 -6x - 8 = 0
x^2 - 3x -4 = 0
(x-4)(x+1) = 0
(x-4)=0 x = 4 three consecutive integers are 4,5,6
(x+1)=0 x = -1 three consecutive integers are -1,0,1
CHECKING our Answers***
4 * 15 = 2*30 = 60
4 * 0 = 2 * 0*1 = 0